OpenStudy (anonymous):

The sum of two positive number is 8. Find the two numbers such that the sum of the cube of one number and the square of the other is a minimum.

OpenStudy (anonymous):

\[x+y=8\] \[x^3+y^2\] is to be minimized. first equation tells us \[y=8-x\] so we have \[f(x)=x^3+(8-x)^2\] so minimize. need some calc i think

OpenStudy (anonymous):

\[f(x)=x^3+x^2-16 x+64\] \[f'(x)=3x^2+2x-16\] set equal zero, find the critical points etc you good from here?

OpenStudy (anonymous):

of course a cubic polynomial has no minimum, but here we have the domain (0,8) so there will be no doubt be a minimum in this interval